OpenStudy (anonymous):

Solve the system of linear equations below. Enter your solution as an ordered pair (x,y). Do not include spaces in your answer. 2x + y = 11 y = 1

OpenStudy (anonymous):

I'd appreciate getting help. I truly don't understand this..

OpenStudy (radar):

Well it appears it is almost solved, y = 1. Now substitute the value 1 for y in the 2x + y = 11 2x + (1) = 11 2x = 11 - 1 2x = 10 x = 5 Did you follow it thus far with understanding?

OpenStudy (anonymous):

Oh, I see! I didnt know it was that simple! Thanks for the help!!! (=

OpenStudy (radar):

In ordert to comply with the problem your answer would be shown as follows (5,1) no spaces.

OpenStudy (anonymous):

Again, thanks for the help! I appreciate it!

OpenStudy (radar):

You're welcome, and good luck with your studies.